A piece of copper block of mass 24g at 230 oC is placed in a copper calorimeter of mass 60 oC containing 60g of water at 31 oC .assuming heat losses are negligible,calculate the final steady temperature of mixture.( specific heat capacity of water =4200 JKg^1K^1 specific heat capacity of copper =400JKg^1K^1).
1
Expert's answer
2020-03-16T13:15:34-0400

From the conservation of energy:


"m_cc_c(T_c-T)=M_cc_c(T-T_w)+m_wc_w(T-T_w)"

"m_cc_c(T_c-T)=(M_cc_c+m_wc_w)(T-T_w)"

"(24)(400)(230-T)=((60)(400)+(60)(4200))(T-31)"

"T=37.7\degree C"